# What Makes Good Data for Alignment? A Comprehensive Study of Automatic Data Selection in Instruction Tuning

### v1

@@ 4 Conclusion
In this paper, we thoroughly investigate the question of what makes good data for alignment. Our research encompasses three controlled studies conducted across three dimensions: complexity, quality and diversity. Throughout these studies, we propose new methods for automatic data selection and train our models, Deita, on the selected data samples. Experimental results demonstrate that Deita is able to achieve superior or comparable performance to the state-of-the-art open-source models with 10x training samples. We release our selected data for others to effectively align models more efficiently.
@@ Results:
[…]

### v2

@@ 4 Conclusion
In this paper, we thoroughly investigate the question of what makes good data for alignment. Our research encompasses three controlled studies conducted across three dimensions: complexity, quality and diversity. Throughout these studies, we propose new methods for automatic data selection and train our models, Deita, on the selected data. Experimental results demonstrate that Deita is able to achieve superior or comparable performance to the state-of-the-art open-source models with 10x training samples. We release our selected data to effectively align models more efficiently.
@@ Results:
We compare Deita-LLaMA16K with the random selection baseline and Vicuna-13B-v1.3. In Table 10, we observe that human preferences align closely with GPT-4 scoring evaluations on MT-Bench. Our data selection strategy demonstrates a significant advantage over random selection in terms of human evaluation. Deita-LLaMA16K performs on par with Vicuna-13B-v1.3 in terms of human evaluation with most of the responses being considered tie by the annotators. However, we note that Deita-LLaMA16K is trained with 20x less data than Vicuna-13B-v1.3 which utilizes 125K training samples.